Calculus I, Fall 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 13

1. Section 7.2 Problems 13, 19, 21.

2. Section 7.3 Problems 1, 11, 15, 19, 37, 45, 57.

3. Section 7.3 Derive Table 7.8 using formula for derivatives of inverse functions.